OpenStudy (anonymous):

Trigonometry - Advanced Functions (A reward shall be given for the best response) Use an addition/subtraction formula to determine the exact value of Tan(5π/4)

OpenStudy (anonymous):

\[5\pi/4 * 180/\pi = 225 degrees\] tan(180+45) = \[\frac{ \tan180+\tan45 }{ 1-\tan(180)\tan(45) }\] Like this?

OpenStudy (anonymous):

\[180 degrees = \pi \] 45 degrees = \[\pi/4\] \[\tan(\pi) = \frac{ \sin(\pi) }{ \cos(\pi) }; \sin(\pi) = 0; \tan(\pi)=0\] \[\tan(\pi/4) = \frac{ \sin(\pi/4) }{ \cos(\pi/4) }; \sin(\pi/4) = \cos(\pi/4); \tan(\pi/4) = 1\] \[\frac{ 0+1 }{ 1-0 } = 1\]

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

np